OpenStudy (anonymous):

Solve the equations a) 3x^5 = 1500 b) 2x (-3) = 128x^ (-5)

OpenStudy (nirmalnema):

is that (3x)^5 or 3x^5

OpenStudy (anonymous):

just 3x^5= 1500

OpenStudy (nirmalnema):

x^5 = 500

OpenStudy (nirmalnema):

b) 2x^-3 = 128x^-5 x^-3 / x^-5 = 128/2 x^2 =64 x = 8

OpenStudy (anonymous):

is that with the ( ) ?

OpenStudy (anonymous):

u wrote b) 2x^-3 = 128x^-5 but the -5 was inside a ( )

OpenStudy (anonymous):

mmmh :/

OpenStudy (anonymous):

never mind nirmalnema my mistake, i think im getting crazy hahah

OpenStudy (nirmalnema):

lol..its ok no problem